### v1

@@ 4 Discussion
In Theorem 3.1, we showed that for any finite $n$-player normal-form game $\Gamma$ and any initial strategy profile ${\mathbf{x}}_{1}\in{\mathbf{X}}$, there exists a satisficing path ${\mathbf{x}}_{1},\dots,{\mathbf{x}}_{T}$ of finite length $T=T({\mathbf{x}}_{1})$ terminating at a Nash equilibrium ${\mathbf{x}}_{T}$. From the proof of Theorem 3.1, one makes the following observations.
  1. 1.
    The length of such a path can be uniformly bounded above as $T({\mathbf{x}}_{1})\leq n$.
  2. 2.
    There exists a collection of strategy update functions $\left\{f^{i}_{\Gamma}:{\mathbf{X}}\to\mathcal{X}^{i}\middle|i\in[n]\right\}$ whose joint orbit is the satisficing path described by the proof of Theorem 3.1. That is, $f^{i}_{\Gamma}({\mathbf{x}}_{t})=x^{i}_{t+1}$ for each player $i\in[n]$, every $10\leq t\leq T-$ 1, and every ${\mathbf{x}}_{1}\in{\mathbf{X}}$, where $x^{i}_{t}$ is player $i$’s component of ${\mathbf{x}}_{t}$ in the satisficing path initialized at ${\mathbf{x}}_{1}$.
The proof of Theorem 3.1 is semi-constructive. At each step along the path, we describe how the next strategy profile should be picked (e.g. ${\mathbf{x}}_{t+1}\in{\rm Worse}({\mathbf{x}}_{t})$), but we do not suggest an algorithm for computing it. In at least one place, namely Case 1 where we put ${\mathbf{x}}_{T}:={\mathbf{z}}_{\star}$, the path construction involves moving jointly to a Nash equilibrium in one step. The computational complexity of such a step is prohibitive (Daskalakis et al., 2009), underscoring that ours is an analytical existence result rather than a computational prescription.
Although we have shown that there exists a discrete-time dynamical system on ${\mathbf{X}}$ that converges to Nash equilibrium in $n$ steps and can be characterized by update functions $\{f^{i}_{\Gamma}\}_{i=1}^{n}$, we note that our possibility result does not contradict the impossibility results of (Hart and Mas-Colell, 2003; Babichenko, 2012) or (Milionis et al., 2023). In particular, the functions $\{f^{i}_{\Gamma}\}_{i=1}^{n}$ need not be (and usually will not be) continuous, violating the regularity conditions of Hart and Mas-Colell (2003) and Milionis et al. (2023), and furthermore the functions $\{f^{i}_{\Gamma}\}_{i=1}^{n}$ depend crucially on the game $\Gamma$ in a way that violates the uncoupledness conditions of (Hart and Mas-Colell, 2003) and (Babichenko, 2012).
@@
While this paper dealt with satisficing paths defined using a best responding constraint, the original definition was stated using an $\epsilon$-best responding constraint, according to which a player who was $\epsilon$-best responding was not allowed to switch its strategy. Putting $\epsilon=0$, one recovers the definition used here, but one may also select $\epsilon>0$, which can be desirable to accommodate for estimation error in multi-agent reinforcement learning applications. The added constraint reduces freedom to switch strategies, and thus makes it more challenging to construct paths starting from a given strategy profile. On the other hand, the collection of Nash equilibria is a strict subset of the set of $\epsilon$-Nash equilibria, and one can attempt to guide the process to a different terminal point in a larger set. At this time, it is not clear to us whether the main result of this paper holds for small $\epsilon>0$. It is clear, however, that the proof technique used here will have to be modified, since we have relied on Lemma 3.4, whose proof involved an indifference condition and invoked the fundamental theorem of algebra, and relaxing to $\epsilon>0$ would render such an argument ineffective.
A second interesting question for future work is whether multi-state Markov games with $n>2$ players have the satisficing paths property. The case with $n=2$ was resolved by Yongacoglu et al. (2023), but the proof technique used there did not generalize to $n\geq 3$. By contrast, our proof technique readily accommodates any number of players, but is designed for stateless normal-form games. Our proof used multi-linearity of the expected reward functions $\{R^{i}\}_{i=1}^{n}$, which does not generally hold in the multi-state setting.
[…]
@@ 5 Conclusion
Satisficing paths can be interpreted as a natural generalization of best response paths in which players may experimentally select their next strategy in periods when they fail to best respond to their counterplayers. While (inertial) best response dynamics drive play to equilibrium in certain well-structured classes of games, such as potential games and weakly acyclic games (Fabrikant et al., 2010), the constraint of best responding limits the efficacy of these dynamics in games with cycles in the best response graph (Pangallo et al., 2019). In such games, best response paths leading to equilibrium do not exist, and multi-agent reinforcement learning algorithms designed to produce such paths will not lead to equilibrium.
@@ Appendix A Properties of the auxiliary functions
We begin by discussing the properties of the auxiliary functions $\{F^{i}:i\in[n]\}$, as they are relevant to characterizing best responses. The facts below are well-known. For a reference, see the text of Maschler et al. (2020).
@@
Appendix B Proof of Lemma 3.4
Recall that in the proof of Theorem 3.1, ${\mathbf{x}}_{\star}$ was defined to be some strategy accessible from ${\mathbf{x}}_{k}\in{\mathbf{X}}$ such that all players unsatisfied at ${\mathbf{x}}_{k}$ were satisfied at ${\mathbf{x}}_{\star}$. The statement of Lemma 3.4 was the following.
  • Lemma 3.4 There exists a sequence $\{{\mathbf{y}}\}_{t=1}^{\infty}$, with ${\mathbf{y}}_{t}\in{\rm NoBetter}({\mathbf{x}}_{k})$ for each $t$, such that $\lim_{t\to\infty}{\mathbf{y}}_{t}={\mathbf{x}}_{\star}$.
Proof B.1.
@@
Fixing $\xi>0$ at a sufficiently small value, by the continuity of the functions $\{F^{i}\}_{i\in[n]}$, we have that ${\mathbf{w}}_{\xi}\in{\rm NoBetter}({\mathbf{x}}_{k})$. By the earlier discussion, we have that ${\mathbf{w}}_{\xi}\notin N_{\zeta}({\mathbf{x}}_{\star})$.
@@
Thus, we see that there exists a sequence $\{{\mathbf{y}}_{t}\}_{t=1}^{\infty}$, with ${\mathbf{y}}_{t}\in{\rm NoBetter}({\mathbf{x}}_{k})$ for all $t$, such that $\lim_{t\to\infty}{\mathbf{y}}_{t}={\mathbf{x}}_{\star}$.

### v2

@@ 4 Discussion
In Theorem 1, we showed that for any finite $n$-player normal-form game $\Gamma$ and any initial strategy profile ${\mathbf{x}}_{1}\in{\mathbf{X}}$, there exists a satisficing path ${\mathbf{x}}_{1},\dots,{\mathbf{x}}_{T}$ of finite length $T=T({\mathbf{x}}_{1})$ terminating at a Nash equilibrium ${\mathbf{x}}_{T}$. From the proof of Theorem 1, one makes the following observations. First, the length of such a path can be uniformly bounded above as $T({\mathbf{x}}_{1})\leq n$. Second, there exists a collection of strategy update functions $\left\{f^{i}_{\Gamma}:{\mathbf{X}}\to\mathcal{X}^{i}\middle|i\in[n]\right\}$ whose joint orbit is the satisficing path described by the proof of Theorem 1. That is, $f^{i}_{\Gamma}({\mathbf{x}}_{t})=x^{i}_{t+1}$ for each player $i\in[n]$, every $10\leq t\leq T-$ 1, and every ${\mathbf{x}}_{1}\in{\mathbf{X}}$, where $x^{i}_{t}$ is player $i$’s component of ${\mathbf{x}}_{t}$ in the satisficing path initialized at ${\mathbf{x}}_{1}$.
The proof of Theorem 1 is semi-constructive. At each step along the path, we describe how the next strategy profile should be picked (e.g. ${\mathbf{x}}_{t+1}\in{\rm Worse}({\mathbf{x}}_{t})$), but we do not suggest an algorithm for computing it. In at least one place, namely Case 1 where we put ${\mathbf{x}}_{T}:={\mathbf{z}}_{\star}$, the path construction involves moving jointly to a Nash equilibrium in one step. The computational complexity of such a step is prohibitive [15], underscoring that ours is an analytical existence result rather than a computational prescription.
Although we have shown that there exists a discrete-time dynamical system on ${\mathbf{X}}$ that converges to Nash equilibrium in $n$ steps and can be characterized by update functions $\{f^{i}_{\Gamma}\}_{i=1}^{n}$, we note that our possibility result does not contradict the impossibility results of [22, 2] or [38]. In particular, the functions $\{f^{i}_{\Gamma}\}_{i=1}^{n}$ need not be (and usually will not be) continuous, violating the regularity conditions of [22] and [38], and furthermore the functions $\{f^{i}_{\Gamma}\}_{i=1}^{n}$ depend crucially on the game $\Gamma$ in a way that violates the uncoupledness conditions of [22] and [2].
@@
While this paper dealt with satisficing paths defined using a best responding constraint, the original definition was stated using an $\epsilon$-best responding constraint, according to which a player who was $\epsilon$-best responding was not allowed to switch its strategy. Putting $\epsilon=0$, one recovers the definition used here, but one may also select $\epsilon>0$, which can be desirable to accommodate for estimation error in multi-agent reinforcement learning applications. The added constraint reduces freedom to switch strategies, and thus makes it more challenging to construct paths starting from a given strategy profile. On the other hand, the collection of Nash equilibria is a strict subset of the set of $\epsilon$-Nash equilibria, and one can attempt to guide the process to a different terminal point in a larger set. At this time, it is not clear to us whether the main result of this paper holds for small $\epsilon>0$. It is clear, however, that the proof technique used here will have to be modified, since we have relied on Lemma 1, whose proof involved an indifference condition and invoked the fundamental theorem of algebra, and relaxing to $\epsilon>0$ would render such an argument ineffective.
A second interesting question for future work is whether multi-state Markov games with $n>2$ players have the satisficing paths property. The case with $n=2$ was resolved by [52], but the proof technique used there did not generalize to $n\geq 3$. By contrast, our proof technique readily accommodates any number of players, but is designed for stateless normal-form games. Our proof used multi-linearity of the expected reward functions $\{R^{i}\}_{i=1}^{n}$, which does not generally hold in the multi-state setting.
In this work, satisficing paths were defined in a way that allowed an unsatisfied player $i$ to change its strategy to any strategy in its set $\mathcal{X}^{i}$, without constraint. This is interesting in many problems where the set of strategies can be explicitly and directly parameterized, but may be unrealistic in games where the set of strategies is poorly understood or in which a player can effectively represent only a subset of its strategies $\mathcal{Y}^{i}\subsetneq\mathcal{X}^{i}$. In such games, the question more relevant for algorithm design is whether the game admits satisficing paths to equilibrium within the restricted subset $\mathcal{Y}^{1}\times\cdots\times\mathcal{Y}^{n}$. This point was implicitly appreciated by both [19] and [20] and explicitly noted in [52]. Some negative results were recently established in [50] for games admitting pure strategy Nash equilibrium when randomized action selection was not allowed and the constrained set was given by $\mathcal{Y}^{i}=\mathbb{A}^{i}$, underscoring the importance of the topology of the sets appearing in the proof of Theorem 1.
@@ 5 Conclusion
Satisficing paths can be interpreted as a natural generalization of best response paths in which players may experimentally select their next strategy in periods when they fail to best respond to their counterplayers. While (inertial) best response dynamics drive play to equilibrium in certain well-structured classes of games, such as potential games and weakly acyclic games [16], the constraint of best responding limits the efficacy of these dynamics in games with cycles in the best response graph [43]. In such games, best response paths leading to equilibrium do not exist, and multi-agent reinforcement learning algorithms designed to produce such paths will not lead to equilibrium.
@@
Appendix: Proofs of technical lemmas
We now discuss the properties of the auxiliary functions $\{F^{i}:i\in[n]\}$ that were used in the proof of Theorem 1, and we prove Lemma 1.
We remark that for each player $i\in[n]$, we identify their set of mixed strategies $\mathcal{X}^{i}=\Delta_{\mathbb{A}^{i}}$ with the probability simplex in $\mathbb{R}^{\mathbb{A}^{i}}$. Thus, $\mathcal{X}^{i}$ inherits the Euclidean metric from $\mathbb{R}^{|\mathbb{A}^{i}|}$. Neighbourhoods and limits in $\mathcal{X}^{i}$ (or its subsets) are defined with respect to this metric. Similarly, we inherit a Euclidean metric for ${\mathbf{X}}$. For $\zeta>0$, we let $N_{\zeta}({\mathbf{x}})$ denote the $\zeta$-neighbourhood of the strategy profile ${\mathbf{x}}\in{\mathbf{X}}$.
@@ Appendix A Properties of the auxiliary functions
We begin by discussing the properties of the auxiliary functions $\{F^{i}:i\in[n]\}$, as they are relevant to characterizing best responses. The facts below are well known. For a reference, see the text of [35].
@@
Appendix B Proof of Lemma 1
Recall that in the proof of Theorem 1, ${\mathbf{x}}_{\star}$ was defined to be some strategy accessible from ${\mathbf{x}}_{k}\in{\mathbf{X}}$ such that all players unsatisfied at ${\mathbf{x}}_{k}$ were satisfied at ${\mathbf{x}}_{\star}$. The statement of Lemma 1 was the following.
  • Lemma 1 If ${\rm Worse}({\mathbf{x}}_{k})=\varnothing$, then there exists a sequence $\{{\mathbf{y}}\}_{t=1}^{\infty}$, with ${\mathbf{y}}_{t}\in{\rm NoBetter}({\mathbf{x}}_{k})$ for each $t$, such that $\lim_{t\to\infty}{\mathbf{y}}_{t}={\mathbf{x}}_{\star}.$
Proof.
@@ Appendix B Proof of Lemma 1
Note that we have defined $w^{i}_{\xi}=x^{i}_{k}$ for $i\in{\rm Sat}({\mathbf{x}}_{k})$, which is to say that we change only the strategies of the unsatisfied players, meaning ${\mathbf{w}}_{\xi}\in{\rm Access}({\mathbf{x}}_{k})$. We will show that if $\xi>0$ is sufficiently small, then continuity of the functions $\{F^{i}\}_{i\in[n]}$ guarantees that ${\mathbf{w}}_{\xi}\in{\rm NoBetter}({\mathbf{x}}_{k})$.
Indeed, player $i\in[n]$ is unsatisfied at ${\mathbf{x}}_{k}$ if and only if it fails to best respond, $x^{i}_{k}\notin{\rm BR}^{i}_{0}({\mathbf{x}}^{-i}_{k})$. Using the function $F^{i}$, this is equivalent to $F^{i}(x^{i}_{k},{\mathbf{x}}^{-i}_{k})>0$. For each player $i\in{\rm UnSat}({\mathbf{x}}_{k})$, let $\sigma_{i}>0$ be such that $F^{i}({\mathbf{x}}_{k})\geq\sigma_{i}>0$. Define $\bar{\sigma}=\min\{\sigma_{i}:i\in{\rm UnSat}({\mathbf{x}}_{k})\}$.
The following statement holds by the continuity of the functions $\{F^{i}\}_{i=1}^{n}$: for each player $i\in[n]$, there exists $e_{i}>0$ such that if a strategy profile ${\mathbf{x}}$ belongs to the $e_{i}$ neighbourhood of ${\mathbf{x}}_{k}$ (i.e. ${\mathbf{x}}\in N_{e_{i}}({\mathbf{x}})$), then $|F^{i}({\mathbf{x}})-F^{i}({\mathbf{x}}_{k})|<\bar{\sigma}/2$. Since $F^{i}({\mathbf{x}}_{k})\geq\bar{\sigma}$, it follows that $F^{i}({\mathbf{x}})>\bar{\sigma}/2>0$, and player $i$ is not best responding at ${\mathbf{x}}\in N_{e_{i}}({\mathbf{x}})$.
Let $\bar{e}:=\min\{e_{i}:i\in[n]\}$. By taking $\xi<\bar{e}/(2n)$, one has that ${\mathbf{w}}_{\xi}\in N_{\bar{e}}({\mathbf{x}}_{k})$. From the preceding remarks, one can see that ${\rm UnSat}({\mathbf{x}}_{k})\subseteq{\rm UnSat}({\mathbf{w}}_{\xi})$, since all players who were unsatisfied at ${\mathbf{x}}_{k}$ remain unsatisfied at ${\mathbf{w}}_{\xi}$. Since $w^{j}_{\xi}=x^{j}_{k}$ for any player $j\in{\rm Sat}({\mathbf{x}}_{k})$, one also has that ${\mathbf{w}}_{\xi}\in{\rm Access}({\mathbf{x}}_{k})$. These two parts combine to show that ${\mathbf{w}}_{\xi}\in{\rm NoBetter}({\mathbf{x}}_{k})$.
Fixing $\xi>0$ at a sufficiently small value ($\xi\in(0,\bar{e}/2n)$), the preceding deductions show that ${\mathbf{w}}_{\xi}\in{\rm NoBetter}({\mathbf{x}}_{k})$. By the earlier discussion, we have that ${\mathbf{w}}_{\xi}\notin N_{\zeta}({\mathbf{x}}_{\star})$.
@@
Thus, we see that there exists a sequence $\{{\mathbf{y}}_{t}\}_{t=1}^{\infty}$, with ${\mathbf{y}}_{t}\in{\rm NoBetter}({\mathbf{x}}_{k})$ for all $t$, such that $\lim_{t\to\infty}{\mathbf{y}}_{t}={\mathbf{x}}_{\star}$. ∎
Appendix C Markov games: model and connections to Theorem 1
Markov games are popular model in the field of multi-agent reinforcement learning. Since the model is quite standard, we offer a short description of the fundamental objects and notations, and we then describe connections between Theorem 1 and a possible extension to multi-state Markov games.
[…]
Description of play. Markov games are played across discrete time, indexed by $t\in\mathbb{N}$. At time $t$, the state variable is denoted $SSs_{t}\in\SS$ and each player $i\in[n]$ selects an action $a^{i}_{t}\in\mathbb{A}^{i}$ according to a distribution $\pi^{i}(\cdot|s_{t})$: $a^{i}_{t}\sim\pi^{i}(\cdot|s_{t})$. The transition probability function $SS\pi^{i}\in\mathcal{P}(\mathbb{A}^{i}|\SS)$ is called player $i$’s policy, and we denote player $i$’s set of policies by $SS\Pi^{i}:=\mathcal{P}(\mathbb{A}^{i}|\SS)$. For any time $t\in\mathbb{N}$, the collection of actions $\{a^{i}_{t}\}_{i=1}^{n}$ is mutually conditionally independent given $s_{t}$. Upon selection of the action profile ${\mathbf{a}}_{t}:=(a^{i}_{t})_{i=1}^{n}$, each player $i$ receives the reward $r^{i}(s_{t},{\mathbf{a}}_{t})$, and the state transitions from $s_{t}$ to $s_{t+1}$ according to $s_{t+1}\sim\mathcal{T}(\cdot|s_{t},{\mathbf{a}}_{t})$.
Player $i$’s performance criterion is its expected $\gamma$-discounted return, which depends on the state variable and the collective policy profile $\bm{\pi}:=(\pi^{1},\dots,\pi^{n})$, which we also denote by $(\pi^{i},\bm{\pi}^{-i})$ to isolate player $i$’s policy. We let $\bm{\Pi}:=\Pi^{1}\times\cdots\times\Pi^{n}$ denote the set of policy profiles. For each pair ($SS(\bm{\pi},s)\in\bm{\Pi}\times\SS$, player $i$’s expected $\gamma$-discounted return is given by
$V^{i}(\pi^{i},\bm{\pi}^{-i},s):=\mathbb{E}_{\bm{\pi}}\left[\sum_{t=1}^{\infty}\gamma^{t-1}r^{i}(s_{t},{\mathbf{a}}_{t})\middle|s_{1}=s\right],$
where $\mathbb{E}_{\bm{\pi}}$ denotes that for every $t\geq 1$, we have that $a^{j}_{t}\sim\pi^{j}(\cdot|s_{t})$ for each player $j\in[n]$ and, implicitly, $s_{t+1}\sim\mathcal{T}(\cdot|s_{t},{\mathbf{a}}_{t})$.
Definition 6.
For $\epsilon\geq 0$, a policy $\pi^{i}_{*}\in\Pi^{i}$ is called an $\epsilon$-best response to $\bm{\pi}^{-i}$ if
$SSV^{i}(\pi^{i}_{*},\bm{\pi}^{-i},s)\geq V^{i}(\pi^{i},\bm{\pi}^{-i},s)-\epsilon,\quad\forall\pi^{i}\in\Pi^{i},\quad\forall s\in\SS.$
Definition 7.
For $\epsilon\geq 0$, a policy profile $\bm{\pi}_{*}=(\pi^{i}_{*},\bm{\pi}^{-i}_{*})\in\bm{\Pi}^{i}$ is called a Markov perfect $\epsilon$-equilibrium if, for each player $i\in[n]$, $\pi^{i}_{*}$ is an $\epsilon$-best response to $\bm{\pi}^{-i}_{*}$.
Putting $\epsilon=0$ into the definitions above, we recover the classical definitions of best responding and Markov perfect equilibrium. In analogy to normal-form games, we use ${\rm BR}^{i}_{\epsilon}(\bm{\pi}^{-i})\subseteq\Pi^{i}$ to denote player $i$’s set of $\epsilon$-best-responses to a given counterplayer policy profile $\bm{\pi}^{-i}$.
Remarks on Markov games
As is conventional in the literature on MARL, we focus on policies that are stationary, Markovian, and possibly randomized. That is, we focus on policies for player $i$ that map states $SSs_{t}\in\SS$ to distributions over the agent’s action set $\mathbb{A}^{i}$ and sample each action $a^{i}_{t}$ from that distribution in a time-invariant and history-independent manner. In principle, agents could use policies that depend also on the time index $t$ or on the history of states and actions. However, the bulk of works on MARL consider this simpler class of policies and this is justifiable for several reasons. We refer the reader to [30] for a summary of such justifications.
Markov games generalize both normal-form games (taking the state space $SS\SS$ to be a singleton) and also MDPs (taking the number of players $n=1$). Moreover, when player $i$’s counterplayers follow a stationary policy $\bm{\pi}^{-i}\in\bm{\Pi}^{-i}$, as assumed in this work, player $i$’s stochastic control problem is equivalent to a single-agent MDP (whose problem data depend on $\bm{\pi}^{-i}$). It follows that player $i$’s set of (stationary) best responses to $\bm{\pi}^{-i}$ is non-empty. Furthermore, player $i$’s best response condition can be characterized using the familiar action value (Q-) functions of reinforcement learning theory. We briefly summarize this below.
In addition to the objective criterion $V^{i}(\pi^{i},\bm{\pi}^{-i},s)$, which is called the value function, we may also define the action value function $Q^{i}$ for player $i$ as
$Q^{i}(\pi^{i},\bm{\pi}^{-i},s,a^{i}):=\mathbb{E}_{\bm{\pi}}\left[\sum_{t=1}^{\infty}\gamma^{t-1}r^{i}(s_{t},{\mathbf{a}}_{t})\middle|s_{1}=s,a^{i}_{1}=a^{i}\right],$
[…]
We further define an optimal action value function for player $i$ against $\bm{\pi}^{-i}$, denoted $Q^{*i}_{\bm{\pi}^{-i}}$, as
[…]
For any policy $\bm{\pi}=(\pi^{i},\bm{\pi}^{-i})$, one can express player $i$’s value function using its Q-function and conditional expectations as $V^{i}(\bm{\pi},s)=\sum_{a^{i}}\pi^{i}(a^{i}|s)Q^{i}(\bm{\pi},s,a^{i})$. From this, it follows that
$SS\max_{a^{i}\in\mathbb{A}^{i}}Q^{*i}_{\bm{\pi}^{-i}}(s,a^{i})=\max_{\pi^{i}_{*}\in\Pi^{i}}V^{i}(\pi^{i}_{*},\bm{\pi}^{-i},s),\quad\forall s\in\SS.$
This equality allows us to characterize best responses using a function $f^{i}:\bm{\Pi}\to\mathbb{R}$, analogous to the function $F^{i}$ appearing in the normal-form case. We define $f^{i}(\bm{\pi})$ as
[…]
The functions $\{f^{i}\}_{i=1}^{n}$ defined above possess the three properties we required of the functions $\{F^{i}\}_{i=1}^{n}$ in the proof of Theorem 1: (a) $f^{i}$ is continuous on $\bm{\Pi}$ [52], (b) $f^{i}(\bm{\pi})\geq 0$ for all $\bm{\pi}\in\bm{\pi}$, and (c) $f^{i}(\pi^{i},\bm{\pi}^{-i})=0$ if and only if $\pi^{i}$ is a best response to $\bm{\pi}^{-i}$.
On extending Theorem 1 to Markov games
We now turn our attention to the task of extending Theorem 1 to Markov games. Following the proof of Theorem 1, virtually all steps can be reproduced in the multi-state setting. To begin, one can construct a satisficing path $\bm{\pi}_{1},\bm{\pi}_{2},\dots,\bm{\pi}_{k}$ by growing the set of unsatisfied players at each iteration until either ${\rm UnSat}(\bm{\pi}_{k})=[n]$ or ${\rm Worse}(\bm{\pi}_{k})=\varnothing$. In the latter case, one can consider the subgame involving only the players in ${\rm UnSat}(\bm{\pi}_{k})$ and obtain a Markov perfect equilibrium $\tilde{\bm{\pi}}_{\star}$ for that subgame, which can then be extended to a policy profile $\bm{\pi}_{\star}\in{\rm Access}(\bm{\pi}_{k})$ by putting
$\pi^{i}_{\star}=\begin{cases}\tilde{\pi}^{i}_{\star},&\text{ if }i\in{\rm UnSat}(\bm{\pi}_{k}),\\ \pi^{i}_{k},&\text{ if }i\in{\rm Sat}(\bm{\pi}_{k}).\end{cases}$
To complete the extension of Theorem 1 to Markov games, one must show that this policy $\bm{\pi}_{\star}\in\bm{\Pi}$ is a Markov perfect equilibrium of the $n$-player Markov game. Since the functions $\{f^{i}\}_{i=1}^{n}$ also satisfy the continuity and semi-definiteness properties described in Appendix A, one possible technique for completing this proof involves showing that the policy $\bm{\pi}_{\star}$ is a limit point of the set ${\rm NoBetter}(\bm{\pi}_{k})$. In other words, one possible technique for completing this proof requires extending Lemma 1 to the multi-state case.
Up to this point, analysis of the stateless case and the multi-state case have been conducted perfectly in parallel. However, it is in the extension of Lemma 1 that the presence of a state leads to a discrepancy in the analysis that will necessitate a novel proof technique for the extension of Theorem 1 to Markov games. We elaborate below on this discrepancy.
Normal-form game analysis.
In the context of finite normal-form games, our proof of Lemma 1 in Appendix B involves a proof by contradiction that exploits the explicit form of an indifference condition in the stateless case. In simple terms, if a player $i$ is best responding and placing positive probability on every action, then any two actions offer equal expected payoff. In symbols, we note that $R^{i}(\delta_{a^{i}_{1}},{\mathbf{x}}^{-i})=R^{i}(\delta_{a^{i}_{2}},{\mathbf{x}}^{-i})$ if and only if
$\sum_{{\mathbf{a}}^{-i}\in{\mathbf{A}}^{-i}}\left[r^{i}(a^{i}_{1},{\mathbf{a}}^{-i})-r^{i}(a^{i}_{2},{\mathbf{a}}^{-i})\right]\mathbb{P}_{{\mathbf{x}}^{-i}}({\mathbf{a}}^{-i})$
$=$ $\sum_{{\mathbf{a}}^{-i}\in{\mathbf{A}}^{-i}}\left[r^{i}(a^{i}_{1},{\mathbf{a}}^{-i})-r^{i}(a^{i}_{2},{\mathbf{a}}^{-i})\right]\prod_{j\not=i}\left\{x^{j}(a^{j})\right\}=0.$
For reasons that will be clarified below, we refer to the expressions $\left[r^{i}(a^{i}_{1},{\mathbf{a}}^{-i})-r^{i}(a^{i}_{2},{\mathbf{a}}^{-i})\right]$ as coefficient terms, and we refer to the terms $\mathbb{P}_{{\mathbf{x}}^{-i}}({\mathbf{a}}^{-i})=\prod_{j\not=i}\left\{x^{j}(a^{j})\right\}$ as strategy-dependent terms. We remark that in the case of normal-form games, the coefficient terms above do not depend on the strategy ${\mathbf{x}}^{-i}$.
Our proof of Lemma 1 in Appendix B considered a one-parameter family of strategies parameterized by $\lambda\in[0,1]$. As part of an intricate proof by contradiction, we obtained an indifference condition, (3), for a player $i^{\dagger}$ who played each action with positive probability while also best responding. Due to the explicit parameterization by $\lambda$ of the strategy ${\mathbf{z}}_{\lambda}$, we are able to recognize that the indifference condition in (3) is characterized by the roots of a polynomial in $\lambda$. Critically, the lefthand-side of (3) is a polynomial in $\lambda$ because the coefficient terms do not depend on the strategy ${\mathbf{z}}_{\lambda}$ and hence do not depend on $\lambda$, while the strategy-dependent terms are polynomials in $\lambda$.
Markov game analysis.
By contrast, we now study indifference conditions in Markov games. Consider an agent $i$ who is best responding to a policy $\bm{\pi}^{-i}$ and places positive probability on actions $a^{i}_{1}$ and $a^{i}_{2}$ in state $s$. The optimality condition is turned into an indifference condition between $a^{i}_{1}$ and $a^{i}_{2}$ in state $s$ as follows:
$Q^{*i}_{\bm{\pi}^{-i}}(s,a^{i}_{1})=Q^{*i}_{\bm{\pi}^{-i}}(s,a^{i}_{2})=\max_{a^{i}\in\mathbb{A}^{i}}Q^{*i}_{\bm{\pi}^{-i}}(s,a^{i}).$
[…]
where $\mathbb{P}_{\bm{\pi}}({\mathbf{a}}^{-i}|s)=\prod_{j\not=i}\pi^{j}(a^{j}|s)$ denotes the probability of the action profile ${\mathbf{a}}^{-i}$ in state $s$ under policy $\bm{\pi}$. In analogy to the normal-form case, we refer to $\mathbb{P}_{\bm{\pi}}({\mathbf{a}}^{-i}|s)$ as the strategy-dependent term and we refer to the term enclosed in square brackets as the coefficient term. However, unlike the normal-form case, here it is clear that the (so-called) coefficient term also depends on the policy $\bm{\pi}^{-i}$, through the term $\max_{a^{i}_{\star}\in\mathbb{A}^{i}}Q^{*i}_{\bm{\pi}^{-i}}(s,a^{i}_{\star})$.
Suppose now that we obtain a one-parameter family of policies $\{\bm{\varpi}_{\lambda}:0\leq\lambda\leq 1\}$ parameterized by some $\lambda\in[0,1]$, in analogy to our construction of ${\mathbf{z}}_{\lambda}$ in Appendix B. Since the coefficient term depends on the policy of player $i$’s counterplayers, one has that the indifference condition
$Q^{*i}_{\bm{\varpi}^{-i}_{\lambda}}(s,a^{i}_{1})-Q^{*i}_{\bm{\varpi}^{-i}_{\lambda}}(s,a^{i}_{2})=0$
cannot generally be characterized by the roots of a polynomial in the parameter $\lambda$ [^4] Although this indifference condition does not generally yield a polynomial in $\lambda$, one can easily find special cases of Markov games in which it does. For instance, if player $i$’s action does not influence transition probabilities, the indifference condition will yield a polynomial and the normal-form proof technique will go through without modification.
Without characterization of the indifference condition as a polynomial in the policy parameter, our proof technique in Appendix B becomes unsuitable for the multi-state setting: we cannot invoke the fundamental theorem of algebra to conclude that the coefficient terms are identically zero, and thus we cannot obtain the contradiction critical to our proof by contradiction, where we found that player $i^{\dagger}$ is in fact indifferent even at the extreme parameter value of $\lambda=1$.
In summary, the proof technique employed in Appendix B to prove Lemma 1 relies crucially on the specific explicit form of the indifference condition in stateless, finite normal-form games. Passing to the multi-state setting, the analogous indifference condition takes a different form, and so the specific derivations cannot be repurposed for a simple extension of Lemma 1. However, it is also important to recognize that this phenomenon is a limitation of the proof technique and does not pose a fundamental obstacle to the generalization of Theorem 1 per se. Indeed, the remaining elements of the proof of Theorem 1 carry over seemlessly to the multi-state case, including various continuity conditions for functions characterizing best responses. It therefore seems promising that one can generalize Theorem 1 to apply to Markov games by applying similar machinery as used in this paper but substituting a different proof for that of Lemma 1 to take advantage of topological or geometric structure shared by both normal-form and Markov games. We leave this as an interesting open question for future research.